OpenStudy (anonymous):

Solve. 5x2 - 3x = (5x2 - 7x) + (8x - 2)

OpenStudy (anonymous):

are those squared

OpenStudy (anonymous):

yeah 5x^2-3x=(5x^2-7x)+(8x-2)

OpenStudy (anonymous):

combine all like terms and isolate x

OpenStudy (anonymous):

\[5x^2-3x=5x^2-7x+8x-2\] Combine like terms \[5x^2-3x=5x^2+x-2\] Now simplify terms \[-4x=-2\] Solve for x \[\frac{-4x}{-4}=\frac{-2}{-4}=\frac{1}{2}\]

OpenStudy (anonymous):

thank u!